OpenStudy (anonymous):

The equation of a linear relation is y = 5x + 9. What is the value of x when y = 34?

OpenStudy (anonymous):

x=5

OpenStudy (jacobbenvenutty):

34 = 5x + 9 -9 - 9 25 = 5x /5 /5 5 = x

OpenStudy (jacobbenvenutty):

Heaven is right

OpenStudy (jacobbenvenutty):

I dont know why she put 54 times 5 though?

OpenStudy (anonymous):

so first you plug in 34 where the y once was; so y=5x+9 will now become 34=5x+9 subtract nine from both sides of the equation 34-9=5x (the nines cancelled out on the right side of the equation) 25=5x divide both sides by five x=25/5 x=?

OpenStudy (anonymous):

do you get the process? @42cookiethere

OpenStudy (anonymous):

@42cookiethere you still there?

OpenStudy (anonymous):

yeah I'm here, this just doesn't really make sense, haha, Answer is...?

OpenStudy (jacobbenvenutty):

Did you not read the answers we put?

OpenStudy (jacobbenvenutty):

If you didn't understand my answer since it is numbers only, read @icecreamman answer. He has words explaining each step

OpenStudy (anonymous):

so you really dont know what 25/5 is?!?!?!